# Recursive subhomogeneity of orbit-breaking subalgebras of $\mathrm{C}^{*}$-algebras associated to minimal homeomorphisms twisted by line bundles

### v1

@@ Abstract.
In this paper, we construct a recursive subhomogeneous decomposition for the Cuntz–Pimsner algebras obtained from breaking the orbit of a minimal Hilbert $C(X)$-bimodule at a subset $Y\subset X$ with non-empty interior. This generalizes the known recursive subhomogeneous decomposition for orbit-breaking subalgebras of crossed products by minimal homeomorphisms.
@@ 1. Introduction
Introduced by Phillips in [32], the similarity of recursive subhomogeneous algebras to direct sums of homogeneous $\mathrm{C}^{*}$-algebras allowed for the extension of constructions required for the study of $K$-theory, stable, real, and decomposition rank, and simple inductive limits of interest to the Elliott classification program [32, 31, 42]. Indeed, when a simple unital $\mathrm{C}^{*}$-algebra $A$ can be modelled as an inductive limit of separable recursive subhomogeneous $\mathrm{C}^{*}$-algebra with slow dimension growth then it is classifiable [13]. In fact, inductive limits of RSH algebras are known to exhaust the Elliott invariant [12].
@@
If $A\cong C(X)$ for a compact metrizable space $X$, then any Hilbert bimodule over $A$ which is finitely generated projective as a right Hilbert $A$-module and is left full, is necessarily given by a right Hilbert module of sections of a line bundle over $X$ with left action given by the right action of $C(X)$ after composing with a homeomorphism $\alpha:X\to X$ [2, Proposition 3.7]. We say that the homeomorphism $\alpha$ is twisted by the line bundle $\mathscr{V}$ and denote such a module by $\Gamma(\mathscr{V},\alpha)$ and the corresponding Cuntz–Pimsner algebra as $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$. (Since $\Gamma(\mathscr{V},\alpha)$ is a Hilbert bimodule, one could equivalently construct the crossed product $C(X)\rtimes_{\Gamma(\mathscr{V},\alpha)}\mathbb{Z}$ in the sense of [1]. We use the Cuntz–Pimsner notation because it is slightly more concise.) When the vector bundle $\mathscr{V}$ is trivial, then $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ is precisely the crossed product of $C(X)$ by the homeomorphism $\alpha$, $C(X)\rtimes_{\alpha}\mathbb{Z}$. Now let $Y\subset X$ be a closed non-empty subset. To this additional data we can associate the orbit-breaking Hilbert $C(X)$-module, $C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha)$, which is the Hilbert subbimodule of $\Gamma(\mathscr{V},\alpha)$ where we have restricted the left multiplication to those functions in $C(X)$ which vanish on the set $Y$. We call the Cuntz–Pimsner algebra $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ an orbit-breaking algebra. When the vector bundle is trivial, $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is isomorphic to an orbit-breaking subalgebra of the crossed product $C(X)\rtimes_{\alpha}\mathbb{Z}$.
@@
From their introduction by Putnam in the case of Cantor minimal systems [34], orbit-breaking algebras have played a critical role in understanding properties and structures of crossed products such as $\mathcal{Z}$-stability and stable rank [26, 40, 41, 36, 35, 43, 25, 3]. They are also interesting $\mathrm{C}^{*}$-algebras in their own right, as sufficient conditions are known for when they are simple and $\mathcal{Z}$-stable. Here, $\mathcal{Z}$-stability refers to the property of tensorial absorption of the Jiang–Su algebra $\mathcal{Z}$—a simple, separable, unital, nuclear, projectionless, infinite-dimensional $\mathrm{C}^{*}$-algebra with unique tracial state and with the same $K$-theory as $\mathbb{Z}$ [21]. That is, a $\mathrm{C}^{*}$-algebra $A$ is $\mathcal{Z}$-stable if $A\otimes\mathcal{Z}\cong A$. As such, they provide dynamical models of $\mathrm{C}^{*}$-algebras that are covered by the classification theorem, which says that unital, simple, separable, nuclear $\mathrm{C}^{*}$-algebras which satisfy the Universal Coefficients Theorem (UCT) and are $\mathcal{Z}$-stable, are classified, up to -isomorphism, by $K$-theory and tracial data (see for example [9, 7, 14, 18, 17, 39, 8]). The UCT is, loosely speaking, a technical requirement for classification results that allows one to lift information from $KK$-theory to $K$-theory. All known nuclear $\mathrm{C}^{*}$-algebras satisfy the UCT, but it remains an open problem to determine whether nuclearity implies the UCT (see [6, 44] for a discussion).
@@
We say that a $\mathrm{C}^{*}$-algebra unital, simple, separable, nuclear, $\mathcal{Z}$-stable $\mathrm{C}^{*}$-algebra which satisfies the UCT is classifiable. It was shown in [11] that orbit-breaking subalgebras of crossed products can realise a wide range of $K$-theory—much wider than the crossed products themselves—and hence a wide range of stably finite, classifiable $\mathrm{C}^{*}$-algebras. The introduction of a non-trivial line bundle has the potential to expand this range, possibly exhausting the range of the Elliott invariant in the stably finite case.
@@
In [2], the authors introduced orbit breaking subalgebras of the Cuntz–Pimsner algebra of a minimal homeomorphism $\alpha$ over a compact metrizable space $X$, twisted by a vector bundle $\mathscr{V}$. When $\mathscr{V}$ is a line bundle and the subspace $Y\subset X$ at which the orbit broken is sufficiently small (for example, a finite set of points), it was proved that $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is a centrally large subalgebra of $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ in the sense of Archey and Phillips [4]. This implies that $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is simple and stably finite. Moreover, if the underlying space has finite dimension, it was shown that orbit breaking subalgebras are $\mathcal{Z}$-stable. To show this, one uses results about the Rokhlin dimension of a homeomorphism [38] on a finite-dimensional metric space to first show that the Cuntz–Pimsner algebra itself is $\mathcal{Z}$-stable. Again, appealing to the fact that it is centrally large, this means that $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is also $\mathcal{Z}$-stable.
@@
This implies, via the classification machinery, that $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is an inductive limit of RSH algebras. However, the result doesn’t say anything about the RSH building blocks in the inductive system. Even when $\mathscr{V}$ is trivial, and we are in the setting of crossed products and their orbit-breaking algebras, this result does not recover the approximately RSH structure given by Q. Lin in an unpublished preprint [27] (see [28] for an outline of the construction). It also gives no information on what happens when the subset $Y$ has non-empty interior (in this case the orbit-breaking algebra will not be simple).
@@
The purpose of this paper is to exhibit an RSH decomposition for the Cuntz–Pimsner algebra $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ when $Y\subset X$ is a closed subset with non-empty interior. This implies that when $Y$ is, for example, a single point, $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is approximately RSH.
@@
Such a construction will allow us to investigate further properties of the $\mathrm{C}^{*}$-algebras associated to minimal homeomorphisms twisted by line bundles and their orbit-breaking subalgebras. In particular, a follow-up paper by the authors will show that if $(X,\alpha)$ is a minimal dynamical system with mean dimension zero (a dynamical generalization of covering dimension introduced by Gromov [19] and developed by Lindenstrauss and Weiss [29]), $\mathscr{V}$ is a line bundle, and $Y\subset X$ is sufficiently small, then both $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ and $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ are $\mathcal{Z}$-stable and hence classifiable. In subsequent work, further properties such as $K$-theory and stable rank will be studied.
@@
In Section 5, we explain some properties of orbit-breaking algebras which are the Cuntz–Pimsner algebras of orbit-breaking bimodules. Then after briefly recalling the recursive subhomogeneous structure of orbit-breaking algebras in the case of crossed products by minimal dynamical homeomorphisms, we outline the construction of the RSH decomposition of orbit-breaking algebras.
Endomorphism bundles of vector bundles are the focus of Section 6. Vector bundles constructed from the line bundles in Section 3 are used in this framework to construct the building blocks of the recursive subhomogeneous.
Section 7 establishes an embedding of the orbit-breaking algebras into a direct of section algebras of endomorphism bundles.
The final section is devoted to the RSH decomposition for the range of the embedding obtained in Section 8. This implies that orbit-breaking algebras are recursive subhomogeneous algebras, which is the ultimate goal of this work.
@@ Acknowledgments
The authors would like to thank Gi Hyun Park, Maria Grazia Viola, Magdalena Georgescu and Dawn Archey for helpful discussions.
@@ 2.1. Hilbert bimodules
We are particularly interested in Hilbert $C(X)$-bimodules, but we begin with recalling definition and some results about Hilbert bimodules.
@@
    a right $B$-valued inner product $\langle\cdot,\cdot\rangle_{B}$ giving $\mathcal{E}$ a right Hilbert $B$-module structure, and
@@
    a left $A$-valued inner product ,${}_{A}\langle\cdot,\cdot\rangle$ giving $\mathcal{E}$ a left Hilbert $A$-module structure,
@@
$\xi_{1}\langle\xi_{2},\xi_{3}\rangle_{B}=\,_{A}\!\langle\xi_{1},\xi_{2}\rangle\xi_{3},$
@@
A Hilbert $A$-$B$-bimodule $\mathcal{E}$ is left, respectively right, full if ,$\overline{{}_{A}\langle\mathcal{E},\mathcal{E}\rangle}=A$, respectively $\overline{\langle\mathcal{E},\mathcal{E}\rangle_{B}}=B$. Here $\overline{\langle\mathcal{E},\mathcal{E}\rangle}$ denotes the closed linear span of $\langle\xi_{1},\xi_{2}\rangle$, $\xi_{1},\xi_{2}\in\mathcal{E}$.
Any Hilbert $A$-$B$-bimodule may also be viewed as a special case of an $A$-$B$ $\mathrm{C}^{*}$-correspondence, that is, a right Hilbert $B$-module equipped with a structure map $\varphi:A\to\mathcal{L}(\mathcal{E})$, which is a -homomorphism from $A$ into the $\mathrm{C}^{*}$-algebra of adjointable operators $\mathcal{L}(\mathcal{E})$. The structure map of a Hilbert $A$-$B$-bimodule is just given by left multiplication. Note that not every $A$-$B$ $\mathrm{C}^{*}$-correspondence can be given the structure of a Hilbert $A$-$B$-bimodule.
@@ 2.1.1. Tensor products of Hilbert bimodules
${}_{\mathcal{E}\otimes\mathcal{F}}\langle\xi_{1}\otimes\eta_{1},\xi_{2}\otimes\eta_{2}\rangle={}_{\mathcal{E}}\langle\xi_{1}{}_{\mathcal{F}}\langle\eta_{1},\eta_{2}\rangle,\xi_{2}\rangle,\quad\xi_{1},\xi_{2}\in\mathcal{E},\eta_{1},\eta_{2}\in\mathcal{F},$
@@
$\langle\xi_{1}\otimes\eta_{1},\xi_{2}\otimes\eta_{2}\rangle_{\mathcal{E}\otimes\mathcal{F}}=\langle\eta_{1},\langle\xi_{1},\xi_{2}\rangle_{\mathcal{E}}\eta_{2}\rangle_{\mathcal{F}},\quad\xi_{1},\xi_{2}\in\mathcal{E},\eta_{1},\eta_{2}\in\mathcal{F},$
If $\mathcal{E}$ is a Hilbert $A$-module then we denote

### v2

@@ Abstract.
In this paper, we construct recursive subhomogeneous decompositions for the Cuntz–Pimsner algebras obtained from breaking the orbit of a minimal Hilbert $C(X)$-bimodule at a closed subset $Y\subset X$ with non-empty interior. This generalizes the known recursive subhomogeneous decomposition for orbit-breaking subalgebras of crossed products by minimal homeomorphisms.
@@ 1. Introduction
Introduced by Phillips in [35], the similarity of recursive subhomogeneous algebras to direct sums of homogeneous $\mathrm{C}^{*}$-algebras allowed for the extension of constructions required for the study of $K$-theory, stable, real, and decomposition rank, and simple inductive limits of interest to the Elliott classification program [35, 34, 45]. Indeed, when a simple unital $\mathrm{C}^{*}$-algebra $A$ can be modelled as an inductive limit of separable recursive subhomogeneous $\mathrm{C}^{*}$-algebras with slow dimension growth, then it is classifiable [13]. In fact, inductive limits of RSH algebras are known to exhaust the Elliott invariant [12].
@@
If $A\cong C(X)$ for a compact metrizable space $X$, then any Hilbert bimodule over $A$ which is finitely-generated projective as a right Hilbert $A$-module and is left full, is necessarily given by a right Hilbert module of sections of a line bundle $\mathscr{V}$ over $X$ with left action given by the right action of $C(X)$ after composing with a homeomorphism $\alpha:X\to X$ [2, Proposition 3.7]. We say that the homeomorphism $\alpha$ is twisted by the line bundle $\mathscr{V}$ and denote such a module by $\Gamma(\mathscr{V},\alpha)$ and the corresponding Cuntz–Pimsner algebra as $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$. (Since $\Gamma(\mathscr{V},\alpha)$ is a Hilbert bimodule, one could equivalently construct the crossed product $C(X)\rtimes_{\Gamma(\mathscr{V},\alpha)}\mathbb{Z}$ in the sense of [1]. We use the Cuntz–Pimsner notation because it is slightly more concise.) When the vector bundle $\mathscr{V}$ is trivial, then $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ is precisely the crossed product $C(X)\rtimes_{\alpha}\mathbb{Z}$ of $C(X)$ by the homeomorphism $\alpha$. Now let $Y\subset X$ be a closed non-empty subset. The orbit-breaking Hilbert $C(X)$-bimodule, $C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha)$, is the Hilbert subbimodule of $\Gamma(\mathscr{V},\alpha)$ where we have restricted the left multiplication to those functions in $C(X)$ which vanish on the set $Y$. We call the Cuntz–Pimsner algebra $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ an orbit-breaking algebra. When the vector bundle is trivial, $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is isomorphic to an orbit-breaking subalgebra of the usual crossed product $C(X)\rtimes_{\alpha}\mathbb{Z}$.
@@
From their introduction by Putnam in the case of Cantor minimal systems [37], orbit-breaking algebras have played a critical role in understanding properties and structures of crossed products such as $\mathcal{Z}$-stability and stable rank [29, 43, 44, 39, 38, 47, 28, 3]. They also are interesting $\mathrm{C}^{*}$-algebras in their own right, as sufficient conditions are known for when they are simple and $\mathcal{Z}$-stable. Here, $\mathcal{Z}$-stability refers to the property of tensorial absorption of the Jiang–Su algebra $\mathcal{Z}$—a simple, separable, unital, nuclear, projectionless, infinite-dimensional $\mathrm{C}^{*}$-algebra with unique tracial state and with the same $K$-theory as $\mathbb{C}$ [23]. That is, a $\mathrm{C}^{*}$-algebra $A$ is $\mathcal{Z}$-stable if $A\otimes\mathcal{Z}\cong A$. As such, they provide dynamical models of $\mathrm{C}^{*}$-algebras that are covered by the classification theorem, which says that unital, simple, separable, nuclear $\mathrm{C}^{*}$-algebras which satisfy the Universal Coefficient Theorem (UCT) and are $\mathcal{Z}$-stable, are classified, up to -isomorphism, by $K$-theory and tracial data (see for example [9, 7, 14, 20, 19, 42, 8]). The UCT is, loosely speaking, a technical requirement for classification results that allows one to lift information from $K$-theory to $KK$-theory. All known nuclear $\mathrm{C}^{*}$-algebras satisfy the UCT, but it remains an open problem to determine whether nuclearity implies the UCT (see [6, 48] for a discussion).
@@
We say that a unital, simple, separable, nuclear, $\mathcal{Z}$-stable $\mathrm{C}^{*}$-algebra which satisfies the UCT is classifiable. It was shown in [11] that orbit-breaking subalgebras of crossed products can realize a wide range of $K$-theory—much wider than the crossed products themselves—and hence a wide range of stably finite, classifiable $\mathrm{C}^{*}$-algebras. The introduction of a non-trivial line bundle has the potential to expand this range, possibly exhausting the range of the Elliott invariant in the stably finite case.
@@
In [2], the authors introduced orbit-breaking subalgebras of the Cuntz–Pimsner algebra of a minimal homeomorphism $\alpha$ over a compact metrizable space $X$, twisted by a vector bundle $\mathscr{V}$. When $\mathscr{V}$ is a line bundle and the subspace $Y\subset X$ at which the orbit broken is sufficiently small (for example, a finite set of points), it was proved that $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is a centrally large subalgebra of $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ as defined by Archey and Phillips in [4]. This implies that $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is simple and stably finite. Moreover, if the underlying space has finite covering dimension, it was shown that orbit-breaking subalgebras are $\mathcal{Z}$-stable. To show this, one uses results about the Rokhlin dimension of a homeomorphism [41] on a finite-dimensional metric space to first prove that the Cuntz–Pimsner algebra itself is $\mathcal{Z}$-stable. Appealing to the fact that $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is centrally large, we conclude that is also $\mathcal{Z}$-stable.
@@
This implies, via the classification machinery, that $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is an inductive limit of RSH algebras. However, the result doesn’t say anything about the RSH building blocks in the inductive system. Even when $\mathscr{V}$ is trivial, and we are in the setting of crossed products and their orbit-breaking algebras, this result does not recover the approximately RSH structure given by Q. Lin in an unpublished preprint [30] (see [31] for an outline of the construction). It also gives no information about what happens when the subset $Y$ has non-empty interior (in this case the orbit-breaking algebra will not be simple).
@@
The purpose of this paper is to exhibit an RSH decomposition for the Cuntz–Pimsner algebra $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ when $Y\subset X$ is a closed subset with non-empty interior.
@@
Such a construction will allow us to investigate further properties of the $\mathrm{C}^{*}$-algebras associated to minimal homeomorphisms twisted by line bundles and their orbit-breaking subalgebras. In particular, a follow-up paper by the authors [17] will show that if $(X,\alpha)$ is a minimal dynamical system with mean dimension zero (a dynamical generalization of covering dimension introduced by Gromov [21] and developed by Lindenstrauss and Weiss [32]), $\mathscr{V}$ is a line bundle, and $Y\subset X$ is sufficiently small, then both $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ and $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ are $\mathcal{Z}$-stable and hence classifiable. In subsequent work, further properties such as $K$-theory and stable rank will be studied.
@@
In Section 4, we recall the definition and properties of orbit-breaking algebras, which are the Cuntz–Pimsner algebras of orbit-breaking bimodules.
In Section 5, we briefly recall the recursive subhomogeneous structure of orbit-breaking algebras in the case of crossed products by minimal homeomorphisms, and we outline the construction of the RSH decomposition of orbit-breaking algebras in our setting.
Endomorphism bundles of vector bundles are the focus of Section 6. Vector bundles constructed from the line bundles in Section 3 are used in this framework to construct the building blocks of the recursive subhomogeneous algebras.
Section 7 establishes an embedding of the orbit-breaking algebras into a direct sum of section algebras of endomorphism bundles.
The final section is devoted to the RSH decomposition for the range of the embedding obtained in Section 7. Indeed, we show that $\mathcal{O}(\mathcal{E}_{Y})$ is isomorphic to the iterated pullback
$\Gamma(\mathscr{M}_{1})\oplus_{\Gamma(\mathscr{M}_{2}|_{\overline{Y_{2}}\setminus Y_{2}})}\oplus\dots\oplus_{\Gamma(\mathscr{M}_{K}|_{\overline{Y_{K}}\setminus Y_{K}})}\Gamma(\mathscr{M}_{K}),$
where $\Gamma(\mathscr{M}_{k})$ are section algebras of locally trivial matrix bundles whose bundle structure is induced by the line bundle $\mathscr{V}$. This implies that the orbit-breaking algebras are recursive subhomogeneous algebras and allows us to conclude our main theorem, Theorem 8.8:
Theorem.
Let $X$ be an infinite compact metric space, $\alpha:X\to X$ a minimal homeomorphism, $\mathscr{V}$ a line bundle over $X$ and $Y\subset X$ a closed subset with non-empty interior. Let $r_{1}<\dots<r_{K}$ denote the distinct first return times to $Y$. The orbit-breaking algebra $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ has a recursive subhomogeneous decomposition $B$ such that
  1. (1)
    the length $l$ of $B$ is at least $K$,
  2. (2)
    the base spaces $Z_{1},\dots,Z_{l}$ have dimension bounded by $X$, and there is $l_{0}=0<l_{1}<l_{2}<\dotsm<l_{K}=l$ such that
    $\bigcup_{j=l_{k-1}+1}^{l_{k}}Z_{j}=\overline{\{y\in Y\mid r(y)=r_{k}\}},$
    for every $1\leq k\leq K$,
  3. (3)
    the matrix sizes are exactly $r_{1},\dots,r_{K}$.
From this, we can recover the classification results from [2], see Corollary 8.9. We end with some remarks about how our RSH construction generalizes that of [30], and we discuss how to obtain more general results concerning classification and stable rank when the base space $X$ is not necessarily of finite dimension.
@@ Acknowledgments
The authors would like to thank Gi Hyun Park, Maria Grazia Viola, Magdalena Georgescu and Dawn Archey for helpful discussions. This project was initiated as part of the workshop 18w5168 Women in Operator Algebras at the Banff International Research Station in November 2018.
@@ 2.1. Hilbert bimodules
We are particularly interested in Hilbert $C(X)$-bimodules, but we begin with recalling the definition and relevant results about Hilbert bimodules over arbitrary $\mathrm{C}^{*}$-algebras.
@@
    a right $B$-valued inner product $\langle\cdot,\cdot\rangle_{\mathcal{E}}$ giving $\mathcal{E}$ a right Hilbert $B$-module structure, and
@@
    a left $A$-valued inner product ${}_{\mathcal{E}}\langle\cdot,\cdot\rangle$ giving $\mathcal{E}$ a left Hilbert $A$-module structure,
@@
$\xi_{1}\langle\xi_{2},\xi_{3}\rangle_{\mathcal{E}}=\,_{\mathcal{E}}\!\langle\xi_{1},\xi_{2}\rangle\xi_{3},$
@@
A Hilbert $A$-$B$-bimodule $\mathcal{E}$ is left, respectively right, full if ,¯$\overline{{}_{\mathcal{E}}\langle\mathcal{E},\mathcal{E}\rangle}\ =A$, respectively ,¯$\overline{\langle\mathcal{E},\mathcal{E}\rangle_{\mathcal{E}}}\ =B$. Here $\overline{\langle\mathcal{E},\mathcal{E}\rangle}$ denotes the closed linear span of $\{\langle\xi_{1},\xi_{2}\rangle\mid\xi_{1},\xi_{2}\in\mathcal{E}\}$.
Any Hilbert $A$-$B$-bimodule may also be viewed as a special case of an $A$-$B$ $\mathrm{C}^{*}$-correspondence, that is, a right Hilbert $B$-module equipped with a structure map $\varphi_{\mathcal{E}}:A\to\mathcal{L}(\mathcal{E})$, which is a -homomorphism from $A$ into the $\mathrm{C}^{*}$-algebra of adjointable operators $\mathcal{L}(\mathcal{E})$. The structure map of a Hilbert $A$-$B$-bimodule is just given by left multiplication. Note that not every $A$-$B$ $\mathrm{C}^{*}$-correspondence can be given the structure of a Hilbert $A$-$B$-bimodule.
@@ 2.1.1. Tensor products of Hilbert bimodules
${}_{\mathcal{E}\otimes_{B}\mathcal{F}}\langle\xi_{1}\otimes\eta_{1},\xi_{2}\otimes\eta_{2}\rangle={}_{\mathcal{E}}\langle\xi_{1}{}_{\mathcal{F}}\langle\eta_{1},\eta_{2}\rangle,\xi_{2}\rangle,\quad\xi_{1},\xi_{2}\in\mathcal{E},\eta_{1},\eta_{2}\in\mathcal{F},$
@@
$\langle\xi_{1}\otimes\eta_{1},\xi_{2}\otimes\eta_{2}\rangle_{\mathcal{E}\otimes_{B}\mathcal{F}}=\langle\eta_{1},\langle\xi_{1},\xi_{2}\rangle_{\mathcal{E}}\eta_{2}\rangle_{\mathcal{F}},\quad\xi_{1},\xi_{2}\in\mathcal{E},\eta_{1},\eta_{2}\in\mathcal{F}.$
If $\mathcal{E}$ is a Hilbert $A$-bimodule then we denote